# Evaluating Privacy Perceptions, Experience, and Behavior of Software Development Teams

### v1

@@ 4 Ethics & Limitations
Ethical Considerations This research adheres to our university’s ethical guidelines and was conducted with our Institutional Review Board (IRB)’s approval. All participants agreed to a thorough consent form that included information about the investigators, the risks, benefits, compensation, and confidentiality. All participants were informed about their voluntary participation, maintaining their right to withdraw at any time. No personally identifiable information was collected; and measures were in place to ensure the anonymity, confidentiality, and security of responses. The contact information of all investigators and the IRB team was also included. No participants contacted the investigators or the IRB about the study or the compensation.
Limitations Like most survey studies, our analysis is based on participant self-report data and is affected by self-report bias, recall bias, and social desirability bias. Participants were informed during consent that the survey pertained to privacy. This may introduce priming and self-selection biases. In addition, there is recruitment bias as the Prolific user base may not fully represent the diverse population of individuals in the SDLC. We used multiple screening questions to ensure that recruited participants in fact have experience in software development activities (Section 3.1). We also asked follow-up and write-in questions to make sure the multiple-choice questions are backed up with written facts. To mitigate the potential for survey responses being generated by AI tools like ChatGPT [10], we minimized open-ended questions in favor of multiple-choice formats and carefully scrutinized the write-in responses to remove those that appeared AI-generated. Short responses with typos and errors suggested that our responses were not AI-generated. Despite our efforts, AI-generated responses could affect the study’s outcomes.
We carefully framed our questions to not prompt biased responses. However, we could not avoid one leading question that asks about the confidence in their companies’ privacy and security measures. We aimed to reduce the bias by providing four options instead of a ‘yes’, ‘no’, with the option to not answer. Additionally, the question follows their own definition of privacy, further helping minimize bias. We employ statistical analyses (like the chi-square test) to ensure the broad applicability of our findings. To control for Type I errors in the presence of multiple hypothesis tests, we report our results after employing Bonferroni correction.
@@ 5 Study Analysis Process
Our survey results are organized around our research questions (RQs, see Section 1), with a focus on various areas of privacy within the SDLC and across different roles. Our RQs examine the perceptions held, privacy experience and challenges, and privacy behaviors, while considering the demographic breakdown (see Section 3.4) to provide additional context and to allow for a more nuanced understanding of the data. Our analysis follows a mixed-method approach, encompassing both quantitative and qualitative methodologies.
Qualitative Analysis We evaluate the descriptive and open-ended questions through open coding procedures and iterative processes. For our open-ended questions, e.g., the definition of privacy, the first two authors, independently, classified 50 responses. They then discussed their results and resolved the discrepancies. They continued with the rest of the responses, evaluated the agreements, and resolved the disagreements in another round of discussion. Lastly, a third privacy expert examined the results to ensure their correctness and completeness. For the non-subjective descriptive questions e.g., which PbD strategies they use, one author categorized them and the second author reviewed them, for correctness.
Quantitative Analysis For the questions where our goal is to understand if a correlation exists between the demographics and the privacy-related perception, experience, and knowledge, we conducted statistical analyses. We used Chi-Squared tests to determine whether there is a significant correlation between two categorical variables. For questions where the responses are on a Likert scale, we used the Kruskal-Wallace test. For perception, experience, and behavioral questions, we hypothesize from our RQs that the size of the company, the presence of a CPO or a similar role, the education level, roles, and participants’ location may impact their confidence in privacy/security measures, various privacy practices such as the creation of PIA or privacy policies and their familiarity with PETs, regulations, and usage of forums.
To control Type I errors and avoid false positives, we use Bonferroni correction [54]. Since Bonferroni correction is very conservative and may increase Type II errors, we discuss the results with respect to $alpha=0.05$ as well as the adjusted value (i.e., $\frac{0.05}{24}=0.0021$, for our 24 statistical tests).
@@ 6.1 Survey Demographics
In our main study, we received a total of 362 responses (after filtering - see Section 3.1). 189 participants reside in the US and the rest 173 come from 22 other countries (see Section 1). Table 1 shows a breakdown of participants’ roles, with the majority ($\sim$55%) in SD roles. As shown in Appendix A - Table 10, the majority of participants identify as male. Most are below the age of 45 and completed their BSc. (61.05%), with $\sim$61in Computer Science (CS), Information Technology (IT), Data Science (DS), and Electrical & Computer Engineering (ECE) majors. This percentage includes the answers to the “Others, please specify” option. Among those with a Business degree, 61% are in AD roles such as product manager and 28% are in SD roles. Among those in the “Other” degree category, 48% identified as SD, 19.5% as QA, 21.0% as AD, and 11.5% as ISec. Their company sizes of $<$100 employees and 100+ employees are distributed almost equally.
@@ 6.2.1 Definition of Privacy
One of the key questions in our survey is “How do you define privacy?”. The responses were diverse and showed differing perceptions. Some participants defined privacy in terms of data security highlighting the need to protect user data from unauthorized access. For example, one participant explained that “It involves implementing measures to safeguard sensitive information, such as encryption, access controls, and data anonymization". Others described privacy from a user rights perspective. For example, “I define privacy as the ability to control all that is related to my information and to keep it from reaching someone who is unauthorized". Few responses incorporated legal compliance, with one participant defining privacy as: “This involves being compliant with regulations and ensuring all data is protected with a least-privilege access model with ownership of the different part data sources with assigned data stewards".
To effectively categorize the diverse definitions of privacy, we utilized Solove’s taxonomy [56], a comprehensive framework that breaks down privacy into various categories based on the types of harm of a privacy breach. We chose Solove’s taxonomy for two key reasons: (a) it provides a structured and detailed approach to understanding and analyzing definitions of privacy, which is essential with our wide range of definitions and perspectives; (b) it has been widely recognized and used in prominent privacy research [7, 31, 77, 9]. We followed an open coding procedure to map the provided definitions with the taxonomy, as described in Section 5. Multiple classes for each definition were also possible. Figure 1 shows the mapping. (For a breakdown of Solove’s Taxonomy see Appendix B - Table 11; the ‘Blackmail’ category did not apply to any participant’s definitions.)
@@
Figure 1 shows that the top frequently occurring categories are ‘Disclosure’, ‘Increased Accessibility’, and ‘Insecurity’. This result indicates that most participants either consider the traditional definition of privacy as control over personal information or perceive privacy in terms of security. For ‘Disclosure’, one participant highlights the importance of transparency and clear communication about data collection purposes and user control: “Privacy is the assurance that all data belonging to an individual will be disclosed to others only with that individual’s consent, for uses understood and approved by that individual.” For ‘Increased Accessibility’, a participant, who works with genetic data, underscores the need for controlled access to such information, only granting access if needed: “The users’ ability to define who can access their data and even in that what kind of data can be accessed. As I work in genetic data from patients in my line of work, the clinical information is always controlled access and only researchers working on the particular project can gain access on a need-to-know basis.”
@@ 6.2.2 Confidence in Security and Privacy Measures
We analyzed whether there is a correlation between participants’ confidence in security and privacy measures and their demographic factors, such as the company’s size (H1a), participants’ roles (H1b), education level (H1c), and the presence of CPO or a similar position (H1d) (see Appendix C and Table 12 for more details). As shown in Table 12, with Bonferroni adjustment ($\frac{0.05}{24}=0.0021$), we cannot reject the null hypothesis for H1a, H1b, and H1c ($p-value=0.494,0.654$ and $0.570$); thus, we find no correlation between confidence in security and privacy measures and company’s size, participants’ roles, or education levels. However, with a $p-value=0.0007$ for H1d, we can reject the null hypothesis and say there is a correlation between the presence of a CPO (or similar position) and confidence in privacy and security measures. We further evaluate the impact of a CPO below.
We asked the ISec members specific questions regarding their company’s security and privacy measures. When asked “whether their company conducts security audits for third-party software used in their products”, slightly more than half ($\sim$56%) said ’Yes’ while a large number ($\sim$38%) were ‘Unsure’. This is alarming since research shows a large number of third-party software and libraries include security and privacy vulnerabilities [74, 27, 2]. However, when we asked “whether their company securely manages encryption keys and implements encryption algorithms and access control policies”, more than 70% responded ‘Yes’ – which highlights inconsistencies in privacy practices even among experts.
@@ 6.2.3 Presence of a Chief Privacy Officer (CPO)
We further evaluate the impact of a CPO or other similar roles on the privacy practices of a company. We focus on the AD and SD roles, who are the majority of our participants (i.e., 268 (74%)). We did not include ISec and QA teams to avoid any response bias due to their active privacy role in the company. We asked them “Do you have a Privacy Officer or similar position in your company?”. Table 13 in Appendix D shows the distribution. Interestingly, only slightly more participants responded ‘Yes’ (42.6%) than ‘No’ (38.4%). $\sim$18were ‘Unsure’, which may indicate the lack of proper communication among employees regarding the company’s privacy practices and the purpose of a CPO. 1.1% responded ‘Other’, which included a legal team or a CTO. Among those that said ‘Unsure’, 23% are in AD roles, and 77% are in SD roles which may indicate CPO members communicate more with the management team (i.e., AD).
@@
We further asked the participants “When you have a question about compliance with regulations, what do you do?”. The participants could select more than one option. Appendix D - Table 15 shows the distribution of the responses. About half of the respondents (50.1%) mention they ask lawyers or a CPO, while 23.1% look at the best practices and standards (such as NIST guidelines), and 18.5% use developers’ forums (such as Stack Overflow). Among ‘Other’ sources, they mainly mention ‘search Internet’ or ‘ask a colleague’.
We analyzed whether the existence of a CPO (i.e., access to a legal or privacy expert) could impact the creation of PIA (H2a), the familiarity with PETs (H2b), the number of privacy breaches (H2c), or is influenced by the company size (H2d). Appendix D and Table 14 show the list of the hypotheses and the results of the tests. With Bonferroni correction, our results show that the presence of a CPO correlates with the size of a company ($p-value=<0.00001$). This correlation indicates that larger companies are more likely to have a CPO or a legal/privacy expert to help mitigate privacy risks, which is aligned with findings in [6]. However, with the p-value adjustment, we do not find a correlation between the presence of a CPO (or a similar position) and the creation of a PIA ($p-value=0.1005$) and the use of PETs ($p-value=0.008$). This may not be surprising, especially since a majority of SD roles, who are the main users of PETs and are involved in the PIA creation, are unaware of a CPO role. Our analysis also did not reveal a significant correlation between the presence of a CPO and the number of privacy breaches experienced by the organization ($p-value=0.359$). This suggests that the presence of a CPO, while important and necessary, may not be sufficient to help minimize privacy breaches.
@@ 6.3.1 Creation of a Privacy Impact Assessment
We received 311 responses, where only 43 (14.0%) of them reported that they created a PIA at any point in the SDLC, while a significant proportion (57.2%) reported that they did not (see Appendix E - Figure 5). This indicates a lack of awareness regarding the existence or the need for PIAs (i.e., the PIAs are non-existent or are conducted without their knowledge by the CPO or other teams). We also observed that $\sim$25are unsure about whether a PIA was created, which may highlight a gap in communication within a company about its privacy practices. $\sim$4chose ‘Prefer not to say’.
Among the 43 who created a PIA, 3 of them ($\sim$7%) did not answer the follow-up questions. Our results show that PIAs were created at various stages in the SDLC (see Figure 2), but almost half are at the planning and analysis stages (51.0%). One participant who reported that a PIA was created during the planning stage said, “At the start of development of idea because privacy is more important than all things”. Some participants reported creating a PIA at the start of development, e.g., “We created a Privacy Impact Assessment at the beginning of the software development process. This allowed us to identify potential privacy risks and develop strategies to mitigate them”. Others mentioned during the design phase, or even towards the end of development.
@@
The responses to the question regarding the process for creating the PIA also varied among participants. Some initiate the process by downloading a template and collaborating with internal specialists, while others seek guidance from lawyers, executives, or third-party experts. A common approach involved consulting with professionals, with one participant mentioning that they "outsourced [a] developer that specialises in data privacy and security". Several participants mentioned the involvement of specific roles, such as the Data Chief, IT teams, and privacy protection specialists. The process often involved cross-functional teams. In some cases, senior leadership, such as the CTO, CEO, or owner, played a pivotal role in the assessment.
Lastly, we evaluated the PIA correlation between the company size (H3a) and the participants’ confidence in privacy and security measures (H3b) (see Appendix E). The results show a significant correlation $p-value<0.0001$ for both tests - even after Bonferroni correction.
@@
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Most members of software teams are not familiar with PIA or are unaware of its creation. However, those involved in PIA emphasize the need for its creation in the initial phases of the SDLC in a collaborative process with expertise from various departments and consultants.

### v2

@@ 4 Ethics & Limitations
Ethical Considerations This research adheres to our university’s ethical guidelines and was conducted with our Institutional Review Board (IRB)’s approval. All participants agreed to a thorough consent form that included information about the investigators, the risks, benefits, compensation, and confidentiality. All participants were informed about their voluntary participation, maintaining their right to withdraw at any time. No personally identifiable information was collected, and measures were in place to ensure the anonymity, confidentiality, and security of responses. The contact information of all investigators and the IRB team was also included. No participants contacted the investigators or the IRB about the study or the compensation.
Limitations Like most survey studies, our analysis is based on participant self-report data and is affected by self-report bias, recall bias, and social desirability bias. Participants were informed during consent that the survey pertained to privacy due to our institutions’ IRB requirement. This may introduce priming and self-selection biases. There is also recruitment bias as the Prolific user base may not fully represent the diverse population of SDLC individuals. We used multiple screening questions to ensure that recruited participants have experience in software development activities (Section 3.1). We adopted a conservative process to remove participants for whom we could not verify their SDLC involvement, however, we may have removed a few professionals. We also asked follow-up and write-in questions to ensure the multiple-choice questions were backed up with written facts. To mitigate the potential for survey responses being generated by AI tools like ChatGPT [15], we minimized open-ended questions in favor of multiple-choice formats and carefully scrutinized the write-in responses to remove those that appeared AI-generated. Short responses with typos and errors suggested that our responses were not AI-generated. Despite our efforts, AI-generated responses could affect the study’s outcomes.
We carefully framed our questions so as not to prompt biased responses. However, we could not avoid one leading question that asks about the confidence in their companies’ privacy and security measures. We aimed to reduce the bias by providing four options instead of a ‘yes’ and ‘no’, with the option to not answer. Additionally, the question follows their own definition of privacy, further helping minimize bias. We employ statistical analyses (like the chi-square test [37]) to ensure the broad applicability of our findings. To control for Type I errors in the presence of multiple hypothesis tests, we report our results after employing Bonferroni correction.
@@ 5 Study Analysis Process
Our survey results are organized around our research questions (RQs, see Section 1), focusing on various areas of privacy within the SDLC and across different roles. Our RQs examine the perceptions held, privacy experience and challenges, and privacy behaviors while considering the demographic breakdown (see Section 3.4) to provide additional context and to allow for a more nuanced understanding of the data. Our analysis follows a mixed-method approach, encompassing both quantitative and qualitative methodologies.
Qualitative Analysis We evaluate the descriptive and open-ended questions through open coding procedures and iterative processes. However, in our analysis, we used taxonomies and categories based on the current literature to classify the responses. For the open-ended question regarding the definition of privacy, the first two authors, independently, classified 50 responses based on the taxonomy of privacy introduced by Solove [79] and the examples and hypotheses from [50, 41]. Similarly, for the usage of PETs, we used PETs categories from the literature [76, 22, 62]. The first two authors independently assigned categories for the first 25 responses. They then discussed their results, resolved the discrepancies, and created a guideline (see Appendix K). They continued with the rest of the responses, evaluated the agreements and resolved the disagreements in another round of discussion. Lastly, a third privacy expert examined the results to ensure their correctness and completeness. For the non-subjective descriptive questions e.g., which PbD strategies they use, one author categorized them based on the current literature, (e.g., privacy by design strategies [47], phases and roles in the SDLC [73] for PIAs) and the second author reviewed them for correctness.
Quantitative Analysis For the questions where our goal is to understand if a correlation exists between the demographics and the privacy-related perception, experience, and knowledge, we conducted statistical analyses. We used the Chi-Squared test [37] to determine whether there is a significant correlation between two categorical variables. For questions where the responses are on a Likert scale, we used the Kruskal-Wallace test [14]. For perception, experience, and behavioral questions, we hypothesize from our RQs that the size of the company, the presence of a CPO or a similar role, the education level, roles, and participants’ location may impact their confidence in privacy/security measures, various privacy practices (such as the creation of PIA or privacy policies), and their familiarity with PETs, regulations, and usage of forums. To control Type I errors and avoid false positives, we use Bonferroni correction [75]. Since Bonferroni correction is very conservative and may increase Type II errors, we discuss the results with respect to $alpha=0.05$ as well as the adjusted value (i.e., $\frac{0.05}{24}=0.0021$, for our 24 statistical tests).
@@ 6.1 Survey Demographics
In our main study, we received a total of 362 responses (after filtering - see Section 3.1). 189 participants reside in the US and the other 173 come from 22 other countries (see Section 1). Table 1 shows a breakdown of participants’ roles, with the majority ($\sim$55%) in SD roles. As shown in Appendix B - Table 10, most participants identify as male, are below the age of 45, and have completed their BSc., with $\sim$61in Computer Science (CS), Information Technology (IT), Data Science (DS), and Electrical & Computer Engineering (ECE) majors. This value includes the answers to “Others, please specify”. Among those with a Business degree, 61% are in AD (e.g., product manager), and 28% are in SD roles. Among those in the “Other” degree category, 48% identified as SD, 19.5% as QA, 21.0% as AD, and 11.5% as ISec. The company sizes of $<$100 and 100+ employees are distributed almost equally.
@@ 6.2.1 Definition of Privacy
One of our key questions is, “How do you define privacy?”. The responses were diverse, showing differing perceptions. Some participants defined privacy in terms of data security, highlighting the need to protect user data from unauthorized access. For example, one participant explained that “It involves implementing measures to safeguard sensitive information, such as encryption, access controls, and data anonymization". Others described privacy from a user rights perspective: “I define privacy as the ability to control all that is related to my information and to keep it from reaching someone who is unauthorized". Few responses incorporated legal compliance, with one participant defining privacy as: “This involves being compliant with regulations and ensuring all data is protected with a least-privilege access model with ownership of the different part data sources with assigned data stewards".
To categorize the diverse definitions of privacy, we utilized Solove’s taxonomy [79], that breaks down privacy into various categories based on the types of harm of a privacy breach. We chose Solove’s taxonomy for two key reasons: (a) it provides a structured and detailed approach to understanding and analyzing definitions of privacy, which is essential with our wide range of definitions and perspectives; (b) it has been widely recognized and used in privacy research [8, 46, 100, 10]. We followed an open coding procedure to map the provided definitions with the taxonomy, as described in Section 5. Multiple classes for each definition were also possible. Figure 1 shows the mapping. (For a breakdown of Solove’s Taxonomy see Appendix C - Table 11; the ‘Blackmail’ category did not apply to any participant’s definitions.)
@@
Figure 1 shows that the top frequently occurring categories are ‘Disclosure’, ‘Increased Accessibility’, and ‘Insecurity’. This result indicates that most participants either consider the traditional definition of privacy as control over personal information or perceive privacy in terms of security. For ‘Disclosure’, one participant highlights the importance of transparency and clear communication about data collection purposes and user control: “Privacy is the assurance that all data belonging to an individual will be disclosed to others only with that individual’s consent, for uses understood and approved by that individual.” For ‘Increased Accessibility’, a participant who works with genetic data underscores the need for controlled access to such information, only granting access if needed: “The users’ ability to define who can access their data and even in that what kind of data can be accessed. As I work in genetic data from patients in my line of work, the clinical information is always controlled access and only researchers working on the particular project can gain access on a need-to-know basis.”
@@ 6.2.2 Confidence in Security and Privacy Measures
We analyzed whether there is a correlation between participants’ confidence in security and privacy measures and their demographic factors, such as the company’s size (H1a), participants’ roles (H1b), education level (H1c), and the presence of CPO or a similar position (H1d) (see Appendix D and Table 12 for more details). As shown in Table 12, with Bonferroni adjustment ($\frac{0.05}{24}=0.0021$), we cannot reject the null hypothesis for H1a, H1b, and H1c ($p-value=0.494,0.654$ and $0.570$); thus, we find no correlation between confidence in security and privacy measures and a company’s size, participants’ roles, or education levels. However, with a $p-value=0.0007$ for H1d, we can reject the null hypothesis and say there is a correlation between the presence of a CPO (or similar position) and confidence in privacy and security measures. We further evaluate whether the existence of a CPO could lead to positive privacy outcomes in Subsection 6.2.3.
We asked the ISec members specific questions regarding their company’s security and privacy measures. When asked “whether their company conducts security audits for third-party software used in their products”, slightly more than half ($\sim$56%) said ‘Yes’ while a large number ($\sim$38%) were ‘Unsure’. This is alarming since research shows a large number of third-party software and libraries include security and privacy vulnerabilities [97, 42, 1]. However, when we asked “whether their company securely manages encryption keys and implements encryption algorithms and access control policies”, more than 70% responded ‘Yes’ – which highlights inconsistencies in privacy practices even among experts.
@@ 6.2.3 Presence of a Chief Privacy Officer (CPO)
To evaluate the impact of a CPO or other similar roles on privacy practices, we focus on the AD and SD roles, who are the majority of our participants (i.e., 268 (74%)). We did not include ISec and QA teams to avoid any response bias, due to their active privacy role in the company. We asked them “Do you have a Privacy Officer or similar position in your company?”. Table 13 in Appendix E shows the distribution. Interestingly, only slightly more participants responded ‘Yes’ (42.6%) than ‘No’ (38.4%). $\sim$18were ‘Unsure’, which may indicate the lack of proper communication among employees regarding the company’s privacy practices and the purpose of a CPO. 1.1% responded ‘Other’, which included a legal team or a CTO. Among those that said ‘Unsure’, 23% are in AD and 77% are in SD roles which may indicate CPO members communicate more with the management team (i.e., AD).
@@
We further asked the participants “When you have a question about compliance with regulations, what do you do?”. The participants could select more than one option. Appendix E - Table 15 shows the distribution of the responses. About half of the respondents (50.1%) mention they ask lawyers or a CPO, while 23.1% look at the best practices and standards (such as NIST guidelines), and 18.5% use developers’ forums (such as Stack Overflow). Among ‘Other’ sources, they mainly mention ‘search Internet’ or ‘ask a colleague’.
We analyzed whether the existence of a CPO (i.e., access to a legal or privacy expert) could impact the creation of PIA (H2a), the familiarity with PETs (H2b), the number of privacy breaches (H2c), or is influenced by the company size (H2d). Appendix E and Table 14 show the list of the hypotheses and the results of the tests. With Bonferroni correction, our results show that the presence of a CPO correlates with the size of a company ($p-value<0.00001$). This correlation indicates that larger companies are more likely to have a CPO or a legal/privacy expert to help mitigate privacy risks, which is aligned with findings in [7]. However, with the p-value adjustment, we do not find a correlation between the presence of a CPO (or a similar position) and the creation of a PIA ($p-value=0.1005$) and the use of PETs ($p-value=0.008$). This may not be surprising, especially since a majority of SD roles, who are the main users of PETs and are involved in the PIA creation, are unaware of a CPO role. Our analysis also did not reveal a significant correlation between the presence of a CPO and the number of privacy breaches experienced by the organization ($p-value=0.359$). This suggests that the presence of a CPO, while important and necessary, may not be sufficient to help minimize privacy breaches.
@@ 6.3.1 Creation of a Privacy Impact Assessment
We received 311 responses, where only 43 (14%) of them (where more than half were outside of EU+UK) reported that they created a PIA at any point in the SDLC, while a significant proportion (57.2%) reported that they did not (see Appendix F - Figure 5). This indicates a lack of awareness regarding the existence or the need for PIAs (i.e., the PIAs are non-existent or are conducted without their knowledge by the CPO or other teams). We also observed that $\sim$25are unsure about whether a PIA was created, which may highlight a gap in communication within a company about its privacy practices. $\sim$4chose ‘Prefer not to say’.
Among the 43 who created a PIA, 3 ($\sim$7%) did not answer the follow-up questions. Our results show that PIAs were created at various stages in the SDLC (see Figure 2), but $\sim$51.0are at the planning and analysis stages. One participant who reported that a PIA was created during the planning stage said, “At the start of development of idea because privacy is more important than all things”. Some participants reported creating a PIA at the start of development, e.g., “We created a [PIA] at the beginning of the software development process. This allowed us to identify potential privacy risks and develop strategies to mitigate them”. Others mentioned during the design, or even towards the end of development.
@@
The responses to the question regarding the process for creating the PIA also varied. Some initiate the process by downloading a template and collaborating with internal specialists, while others seek guidance from lawyers, executives, or third-party experts. A common approach involved consulting with professionals, with one participant mentioning that they "outsourced [a] developer that specialises in data privacy and security". Several participants mentioned the involvement of specific roles, such as the Data Chief, IT teams, and privacy protection specialists. The process often involved cross-functional teams. In some cases, senior leadership, e.g., the CTO, CEO, or owner, played a pivotal role in the process.
Lastly, we evaluated the PIA correlation between the company size (H3a) and the participants’ confidence in privacy and security measures (H3b) (see Appendix F). The results show a significant correlation $p-value<0.0001$ for both tests - even after Bonferroni correction.
@@
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Most members of software teams are not familiar with PIA or are unaware of its creation. However, those involved in PIA emphasize the need for its creation in the initial phases of the SDLC in a collaborative process with experts from various departments and consultants.